12 gal. of a saline solution was mixed with 3 gal. of a 30% saline solution to make a 38% saline solution. Find the percent concentration of the first solution.
Answer by fractalier(6550)

Call the unknown percentage, x. We then write
12x + 3(30) = 15(38)
12x + 90 = 570
12x = 480
x = 40%